# Ding Chen1  Qingchen Yu211footnotemark: 1  Pengyuan Wang211footnotemark: 1  Wentao Zhang322footnotemark: 2  &Bo Tang2  Feiyu Xiong2  Xinchi Li1  Minchuan Yang1  Zhiyu Li [^2] footnotemark: 2

### v1

@@ 4.1.2 Dataset Partitioning
The training and test sets are used to train and evaluate the xVerify model. Both are sampled from the same pool, sharing similar distributions. Specifically, they include samples generated by 15 LLMs across 17 evaluation sets, covering the four previously mentioned question types. The training set contains 36,941 samples, and the test set includes 5,194 samples.
@@ 4.1.3 Data Annotations
To ensure the accuracy of xVerify’s training and evaluation, we conducted multiple rounds of automatic and manual annotation across the three datasets. Specifically, we used GPT-4o to perform two rounds of annotation for all samples in the datasets, utilizing two distinct prompt templates (details provided in Appendix D.2) to improve annotation confidence [33, 22]. Given the large size of the training set, we only applied manual annotation to the more challenging math problems and to samples where the two rounds of GPT-4o annotations disagreed. In contrast, for the test and generalization sets, we manually annotated all samples, resulting in a three-round annotation process to maximize label reliability. Details of the manual annotation process are provided in Appendix B.2.
4.1.4 Data Augmentation
For multiple-choice questions, we applied two types of augmentations: option index transformation and noise injection. The former converts alphabetical labels to Arabic or Roman numerals, while the latter randomly adds or removes irrelevant distractor options without changing the original question intent, thereby increasing structural complexity.
Together, these strategies expanded the expressive space of the dataset while preserving semantic consistency, offering richer and more challenging training signals for xVerify. After augmentation, the sizes of the training, test, and generalization sets increased to 43,204, 6,122, and 6,468 samples respectively. Full dataset details are provided in Appendix B.1. The augmentation of math problems primarily relied on GPT-4o; prompt templates are listed in Appendix D.3.
4.2 Model Training
We trained 14 models with different parameter sizes and architectures using the training set from the VAR dataset. Specifically, we utilized the LLaMA-Factory framework [38] and QLoRA technique [4] for model training. Based on extensive experimentation, we set the number of epochs to 1 and selected a learning rate of 1e-4 as the optimal configuration, with other hyperparameters detailed in Appendix C.1. Many researchers have pointed out potential bias in using LLMs as judge models, where models from the same family tend to receive higher ratings [19]. To thoroughly evaluate the generalization capability of the xVerify method, we trained 14 models with varying parameter sizes and architectures. These models ranged from 0.5B to 32B parameters and included five different families, such as LLaMA 3 [6], Qwen2.5 [37], and Gemma 2 [30]. Details of the models used are provided in Appendix C.2.
5 Experiments
In this section, we will present the configuration, results, and detailed analysis of the xVerify model evaluation experiments. First, we will outline the experimental setup:
  • Datasets: The datasets used in the evaluation experiments are the test set and generalization set from the VAR dataset. The test set is used to evaluate the xVerify model’s performance, while the generalization set supplements the test set by simulating real-world scenarios with a broader sample distribution to assess the model’s generalization ability.
  • Metrics: The evaluation mainly uses accuracy and F1 score on both the test and generalization sets. Accuracy shows the model’s overall performance, while the F1 score combines precision and recall for a more complete perspective.
  • Baselines: There are two types of baselines: evaluation frameworks and judge models. The evaluation frameworks include DeepSeek-Math [29], LM Eval Harness [5], Math-Verify [13], OpenAI Evals [25], OpenCompass [27], and UltraEval [9]. The judge models include PandaLM [34], Auto-J [21], Prometheus 2 [17], JudgeLM [39], and CompassJudger [1]. In addition, GPT-4o is also used as a judge model with two strategies: one with CoT and one without. The prompts for the judge model and xVerify are provided in Appendix D.4 and Appendix D.5.
Test Set Evaluation Results.
We evaluated all evaluation frameworks, judge models, and the xVerify model on the VAR test set (see Table 1). Overall, the xVerify model outperforms all evaluation frameworks and judge models, including GPT-4o, with the best and second-best values in each column appearing for the xVerify model.
Table 1: Evaluation Accuracy Results on the Test Set. "-" indicates that the evaluation method is not applicable to the problem type. The best performance in each column will be shown in bold, and the second-best performance will be underlined.
@@ 5 Experiments
Among the evaluation frameworks, the best performers were DeepSeek Math Verify and OpenCompass, but neither achieved an F1 score nor accuracy exceeding 80%. Some evaluation frameworks were also not suitable for certain question types, which is an inherent limitation of rule-based methods—strong in specificity but limited in applicability. For instance, OpenCompass was completely unsuitable for short answer and classification questions. Additionally, the long reasoning processes generated by reasoning models made it difficult for evaluation frameworks to extract final answers, lowering their overall performance.
Among judge models, GPT-4o and CompassJudger showed the best overall performance. The CompassJudger-1-32B model achieved F1 score and accuracy of 91.67% and 91.69%, respectively. However, the model performed poorly on math questions, with both F1 score and accuracy below 85%, indicating that it handles simpler questions well but struggles with formula equivalence in math problems. Furthermore, only the 32B version of this judge model achieved over 90% F1 score and accuracy, while smaller models performed below 80%. Therefore, the performance of CompassJudger-1-32B is more a result of the base model’s capabilities rather than the subsequent training. For example, the smallest xVerify-0.5B-I model outperforms CompassJudger-1-32B across the board, indicating that the VAR training set significantly improves model evaluation performance. GPT-4o’s overall performance is very close to xVerify, but the improvement after using CoT is small, with token consumption nearly doubling. Specifically, GPT-4o as Judge evaluated the entire test set at a cost of \$13.09, while GPT-4o as Judge (CoT) cost \$20.15 (using the OpenAI API, charged by token count).
In contrast, even the smallest xVerify-0.5B-I model outperforms all methods except GPT-4o as Judge (CoT) in overall performance, and the xVerify-3B-Ib model surpasses all others in every evaluation metric. Moreover, for more difficult math questions, all xVerify models except xVerify-0.5B-I exceeded 95% performance. We also found that the performance of the xVerify model improves as the parameter size increases, but slightly decreases after exceeding 7B parameters, likely due to overfitting on the VAR training set, which is sufficiently large for smaller models.
Generalization Set Evaluation Results.
To better assess the performance of xVerify on a broader sample distribution, we evaluated all methods on the VAR generalization set, as shown in Table 2. On the generalization set, the xVerify model showed a slight decrease in overall performance. However, the drop in both F1 score and accuracy was less than 1.5%, while other methods showed mixed results. Overall, the xVerify model still outperformed all other methods, indicating that although overfitting exists in xVerify, it is limited and the model maintains strong generalization ability on samples outside the training set distribution.
@@
Table 2: Evaluation Accuracy Results on the Generalization Set. "-" indicates that the evaluation method is not applicable to the problem type. The best performance in each column will be shown in bold, and the second-best performance will be underlined.
@@
Specifically, the overall F1 score and accuracy of all evaluation frameworks remained below 80%, with only OpenCompass achieving an overall accuracy above 70%. This indicates that rule-based evaluation frameworks have significant limitations in generalization performance, struggling to effectively handle the diverse answers and evaluation sets from LLMs.
Among the judge models, the best-performing ones remained GPT-4o and CompassJudger. However, all judge models except GPT-4o had an F1 score below 75% on math questions, with most models scoring below 50%, indicating that judge models almost entirely fail in evaluating more diverse and complex math problems. GPT-4o as Judge and GPT-4o as Judge (CoT) also failed to achieve an F1 score above 90% on math problems, suggesting that the math samples in the generalization set indeed present challenges for evaluation methods. Furthermore, GPT-4o’s performance did not improve after using CoT; instead, it showed a slight decline. This suggests that in broader scenarios, CoT-based prompt engineering methods do not effectively improve GPT-4o’s performance as a judge model, and model fine-tuning may be a better option.
In contrast, the xVerify-0.5B-I model outperformed all evaluation methods except GPT-4o, and the xVerify-3B-Ib model outperformed both CoT-based GPT-4o methods. For more difficult math problems, the F1 score and accuracy of the xVerify-14B-Ia and xVerify-32B-I models exceeded 90%. Additionally, we observed that as the parameter size of the xVerify model increased, the performance drop on the generalization set decreased. For example, the accuracy drop for xVerify-0.5B-I was 1.33%, 0.91% for xVerify-9B-I, and 0.80% for xVerify-32B-I, suggesting that larger xVerify models exhibit stronger generalization performance.
Furthermore, we comprehensively evaluated the performance of 14 xVerify models on both the test and generalization sets, and tested the computational efficiency of all xVerify and judge models, along with the evaluation cost of GPT-4o as a judge model. The results showed that xVerify models outperform other judge models in both usage cost and evaluation efficiency. Full experimental results can be found in Appendix E.
@@ 6 Conclusion
In this paper, we propose an efficient answer verifier for reasoning model evaluations, named xVerify, which can effectively assess the correctness of long reasoning responses generated by reasoning models on various difficult objective questions. To train and evaluate the xVerify model, we constructed the VAR dataset based on several popular LLMs and evaluation sets. This dataset primarily collects long reasoning responses generated by reasoning models on challenging questions, and multiple rounds of labeling and verification were conducted using GPT-4o and human annotators. Ultimately, we trained multiple xVerify models of varying specifications based on the VAR dataset and performed comparative evaluations with several evaluation frameworks and judge models on both the test and generalization sets. The experimental results show that even the smallest xVerify-0.5B-I model outperforms all methods except GPT-4o, and larger xVerify models surpass all other methods, demonstrating the effectiveness and generalization ability of xVerify.
@@ Appendix A Datasets and Models
In this study, we employ a total of 24 datasets, which are categorized into four primary types: multiple-choice questions (Choice), short answer questions (Short Answer), mathematical problems (Math), and classification tasks (Classification), as summarized in Table  3. To evaluate the multilingual capabilities of the xVerify model, each question type includes datasets in both Chinese and English, with one dataset featuring multilingual content. For each dataset, samples are partitioned into training and test sets following a 2:1 ratio, with the training and test sets ideally comprising 2,000 and 1,000 instances, respectively. In certain cases, the number of available samples is below 3,000, or the official test set is not publicly available, resulting in reduced dataset sizes after preprocessing.
@@
CMMLU Choice 2000 1000 Chinese CC-BY-NC-4.0
C-Eval Choice 1346 260 Chinese CC-BY-NC-SA-4.0
GPQA Choice 794 398 English CC-BY-4.0
@@
MMLU-Redux Choice 2000 1000 English CC-BY-4.0
@@
CMATH Math 1128 565 Chinese CC-BY-4.0
@@
LiveMathBench Math 190 93 English & Chinese CC-BY-4.0
@@
MGSM Math 1892 946 Multilingual CC-BY-SA-4.0
@@
ARC Short Answer 2000 1000 English CC-BY-SA-4.0
@@
C-SimpleQA Short Answer 2000 1000 Chinese CC-BY-NC-SA-4.0
DROP Short Answer 2000 1000 English CC-BY-SA-4.0
@@
A total of 19 large language models (LLMs) are utilized in our experiments, encompassing a diverse range of model sizes and types, with a particular emphasis on reasoning models (see Table  4). These models are subsequently used to collect LLM-generated responses and to train the xVerify model.
@@ B.1.1 Training Set
The training set comprises 43,204 samples. Tables 5 to 8 provide the sample counts corresponding to each LLM, dataset, prompt template, and question type. Note that datasets with names containing "_enh" refer to the augmented multiple choice question datasets.
Table 5: Number of samples from each LLM in the training set.
@@
Table 6: Number of samples from each dataset in the training set.
@@
Table 8: Number of samples from each question type in the training set.
@@ B.1.2 Test Set
The test set comprises 6,122 samples. Tables 9 to 12 provide the sample counts corresponding to each LLM, dataset, prompt template, and question type. Note that datasets with names containing "_enh" refer to the augmented multiple choice question datasets.
Table 9: Number of samples from each LLM in the test set.
@@
Table 10: Number of samples from each dataset in the test set.
@@
Table 11: Number of samples from each prompt template in the test set.
@@
Table 12: Number of samples from each question type in the test set.
@@ B.1.3 Generalization Set
The generalization set comprises 6,468 samples. Tables 13 to 16 provide the sample counts corresponding to each LLM, dataset, prompt template, and question type. Note that datasets with names containing "_enh" refer to the augmented multiple choice question datasets.
Table 13: Number of samples from each LLM in the generalization set.

### v2

@@ 4.1.2 Dataset Partitioning
The training and test sets are used to train and evaluate xVerify. Both are sampled from the same pool, sharing similar distributions. Specifically, they include samples generated by 15 LLMs across 17 evaluation sets, covering the four question types. The training set contains 36,941 samples, and the test set includes 5,194 samples.
@@ 4.1.3 Data Annotations
To ensure the accuracy of xVerify’s training and evaluation, we conducted multiple rounds of automatic and manual annotation across the three datasets. Specifically, we used GPT-4o to perform two rounds of annotation for all samples in the datasets, utilizing two distinct prompt templates (details provided in Appendix E.2) to improve annotation confidence (self_consistency_2023_iclr_google; self_feedback_2024_arXiv_iaar). Given the large size of the training set, we only applied manual annotation to the more challenging math problems and to samples where the two rounds of GPT-4o annotations disagreed. In contrast, for the test and generalization sets, we manually annotated all samples, resulting in a three-round annotation process to maximize label reliability. Details of the manual annotation process are provided in Appendix B.2.
@@
Table 1: Evaluation Results on the Test Set. "-" indicates that the evaluation method is not applicable to the problem type. Best and second-best results are bold and underlined, respectively
4.1.4 Data Augmentation
For multiple-choice questions, we applied two augmentations: option index transformation and noise injection. The former converts alphabetical labels to Arabic or Roman numerals, while the latter randomly adds or removes irrelevant distractor options without changing the original question intent, thereby increasing structural complexity.
Together, these strategies expanded the expressive space of the dataset while preserving semantic consistency, offering richer and more challenging training signals for xVerify. After augmentation, the sizes of the training, test, and generalization sets increased to 43,204, 6,122, and 6,468 samples respectively. Full dataset details are provided in Appendix B.1. The augmentation of math problems primarily relied on GPT-4o; prompt templates are listed in Appendix E.3.
4.2 Model Training
We trained 14 models with different parameter sizes and architectures using the training set from the VAR dataset. Specifically, we utilized the LLaMA-Factory framework (llamafactory_2024_acl_buaa) and QLoRA technique (dettmers2023qlora) for model training. Based on extensive experimentation, we set the number of epochs to 1 and selected a learning rate of 1e-4 as the optimal configuration, with other hyperparameters detailed in Appendix C.1. Many researchers have pointed out potential bias in using LLMs as judge models, where models from the same family tend to receive higher ratings (leakage_2025_arXiv_arizona). To thoroughly evaluate the generalization capability of the xVerify method, we trained 14 models with varying parameter sizes and architectures. These models ranged from 0.5B to 32B parameters and included five different families, such as LLaMA 3 (llama3_2024_arXiv_meta), Qwen2.5 (qwen25_2024_arXiv_qwen), and Gemma 2 (gemma2_2024_arXiv_google). Details of the models used are provided in Appendix C.2.
5 Experiments
This section details our experiments, covering two main experiments: evaluating xVerify as a judge model on both in-domain and out-of-distribution datasets, and using xVerify as a reward model in the reinforcement learning optimization process. First, we will outline the experimental setup:
Datasets:
For the evaluation experiments, we primarily use the test set and generalization set from the VAR dataset. The test set evaluates xVerify’s core performance, while the generalization set assesses its robustness on out-of-distribution samples. For the reinforcement learning experiments, the dataset for training and testing the policy model is collected from diverse sources. Further details are provided in Appendix D.2.
Metrics:
Accuracy and F1 are used as the main metrics to provide a comprehensive assessment.
Baselines:
There are two types of baselines: evaluation frameworks and judge models. The evaluation frameworks include DeepSeek-Math (deepseekmath_2024_arXiv_deepseek), LM Eval Harness (lm_eval_harness_21_Zenodo), Math-Verify (math_verify_24_github), OpenAI Evals (evals_24_github_OPENAI), OpenCompass (opencompass_24_github), and UltraEval (ultraeval_24_Tsinghua). The judge models include PandaLM (pandalm_24_icrl_Peking), Auto-J (auto_j_24_iclr_JiaoTong), Prometheus 2 (prometheus_2_24_acl_KAIST), JudgeLM (judgelm_25_icrl), and CompassJudger (CompassJudger_1_24_arXiv_Shanghai_AI). GPT-4o is also used as a judge model, with and without CoT. All prompts employed for both the judge model and xVerify can be found in Appendix E.4 and E.5.
@@ 5 Experiments
Table 2: Evaluation Results on the Generalization Set. "-" indicates that the evaluation method is not applicable to the problem type. Best and second-best results are bold and underlined, respectively
Model MMLU-Pro GPQA MATH-500 DROP Amazon CMMLU CHID Avg.
Qwen2.5-7B
Direct Generation 40.4% 18.8% 39.6% 70.4% 94.0% 76.8% 42.4% 54.6%
   $\hookrightarrow$ RL with Math Verify 52.0% 41.2% 79.5% 84.8% 96.0% 88.0% 64.0% 72.2%
   $\hookrightarrow$ RL with xVerify 53.6% 39.6% 81.5% 86.0% 98.4% 88.4% 63.6% 73.0%
Llama3.1-8B
Direct Generation 42.8% 28.8% 52.6% 82.4% 96.0% 55.2% 42.4% 57.2%
   $\hookrightarrow$ RL with Math Verify 46.4% 32.4% 57.9% 88.8% 97.2% 58.0% 42.4% 60.4%
   $\hookrightarrow$ RL with xVerify 46.4% 34.4% 58.0% 87.6% 97.6% 59.6% 44.8% 61.2%
Table 3: Evaluation Accuracy Results of RL with xVerify as Reward Model.
5.1 Evaluation with xVerify as Judge Model
We evaluated all frameworks and models on VAR’s test and generalization sets (Tables 12), where xVerify consistently achieved the best results.
Test Set Evaluation Results.
On the VAR test set, xVerify consistently outperforms all evaluation frameworks and judge models. Even the smallest xVerify-0.5B-I achieves second-best overall accuracy (96.85%) and F1 (96.69%), surpassing CompassJudger-1-32B and matching GPT-4o’s performance while using far fewer tokens. Larger xVerify variants (3B–32B) further improve both F1 and accuracy, peaking at 97.50%/97.41% (F1/Acc.) with xVerify-7B-I. Notably, all xVerify models above 0.5B exceed 95% on challenging math questions, and performance gains taper beyond 7B parameters—suggesting a sweet spot around mid-scale models for this dataset.
Generalization Set Evaluation Results.
On the VAR generalization set, xVerify’s overall F1 and accuracy drop by less than 1.5%, demonstrating strong robustness to out-of-distribution samples. Even xVerify-0.5B-I retains 95.53% accuracy, outperforming all rule-based frameworks and most judge models except GPT-4o. Larger xVerify models reduce the performance gap further: xVerify-14B-Ia reaches 96.65% accuracy with over 90% on math questions. These results confirm that scaling xVerify enhances generalization, and that fine-tuned judge models can outperform CoT-based prompting without incurring extra token costs.
Supplementary experiments detailed in Appendix F provide further empirical support for the effectiveness and robustness of xVerify.
5.2 Reinforcement Learning with xVerify as Reward Model
We investigate whether a more accurate reward model improves optimization efficiency and final performance in RL fine-tuning. We build on veRL (sheng2024hybridflow) and train Qwen2.5-7B and Llama3.1-8B with GRPO (deepseekmath_2024_arXiv_deepseek), using xVerify-7B-I as the reward and Math-Verify as a rule-based baseline. This experiment is a proof-of-concept designed to compare reward-signal quality and training dynamics rather than to maximize final scores. Training hyperparameters are listed in Appendix D.1. For faithful evaluation, we manually annotate all samples (Appendix B.2).
[…]
Furthermore, We show that xVerify aligns closely with human judgments, with full results provided in Appendix F.3.
@@ 6 Conclusion
In this paper, we propose xVerify, an efficient answer verifier for evaluating long reasoning responses generated by reasoning models on challenging objective questions. To train and evaluate xVerify, we construct the VAR dataset based on multiple LLMs and evaluation benchmarks, which collects long-form reasoning answers and is carefully annotated through multiple rounds of verification by GPT-4o and human annotators. Using VAR, we train xVerify models of different scales and compare them with existing evaluation frameworks and judge models on both test and generalization sets. Experimental results show that even the smallest xVerify-0.5B-I model outperforms all methods except GPT-4o, while larger xVerify models achieve the best overall performance, demonstrating strong effectiveness and generalization. Additionally, reinforcement learning experiments show that xVerify is effective as a reward model, effectively enhances policy performance compared to direct generation.
Limitations
In this work, we focus on building a more accurate and efficient answer equivalence verifier, particularly for assessing the equivalence between the outputs of reasoning models and reference answers. However, the current xVerify model is not yet capable of fully replacing all evaluation methods and reward models. On the one hand, xVerify is currently tailored to objective questions with definitive answers and lacks effective optimization for subjective questions. On the other hand, xVerify cannot yet serve as a complete substitute for general-purpose reward models, as such models typically require only the input question and the output from the policy model to generate reward signals, whereas xVerify still depends on the availability of reference answers. Addressing these limitations may require future efforts to enhance xVerify’s capability in scenarios without reference answers, such as those involving subjective questions. This could involve the development of new task settings and corresponding datasets for training and optimizing the xVerify model.
[…]
Appendix
@@ Appendix A Datasets and Models
In this study, we employ a total of 24 datasets, which are categorized into four primary types: multiple-choice questions (Choice), short answer questions (Short Answer), mathematical problems (Math), and classification tasks (Classification), as summarized in Table  4. To evaluate the multilingual capabilities of the xVerify model, each question type includes datasets in both Chinese and English, with one dataset featuring multilingual content. For each dataset, samples are partitioned into training and test sets following a 2:1 ratio, with the training and test sets ideally comprising 2,000 and 1,000 instances, respectively. In certain cases, the number of available samples is below 3,000, or the official test set is not publicly available, resulting in reduced dataset sizes after preprocessing.
@@
CMMLU Choice 2000 1000 Chinese $C$-BY-NC-4.0
C-Eval Choice 1346 260 Chinese $C$-BY-NC-SA-4.0
GPQA Choice 794 398 English $C$-BY-4.0
@@
MMLU-Redux Choice 2000 1000 English $C$-BY-4.0
@@
CMATH Math 1128 565 Chinese $C$-BY-4.0
@@
LiveMathBench Math 190 93 English & Chinese $C$-BY-4.0
@@
MGSM Math 1892 946 Multilingual $C$-BY-SA-4.0
@@
ARC Short Answer 2000 1000 English $C$-BY-SA-4.0
@@
C-SimpleQA Short Answer 2000 1000 Chinese $C$-BY-NC-SA-4.0
DROP Short Answer 2000 1000 English $C$-BY-SA-4.0
@@
A total of 19 large language models (LLMs) are utilized in our experiments, encompassing a diverse range of model sizes and types, with a particular emphasis on reasoning models (see Table  5). These models are subsequently used to collect LLM-generated responses and to train the xVerify model.
@@ B.1.1 Training Set
The training set comprises 43,204 samples. Tables 6 to 9 provide the sample counts corresponding to each LLM, dataset, prompt template, and question type. Note that datasets with names containing "_enh" refer to the augmented multiple choice question datasets.
@@
Table 6: Number of samples from each LLM in the training set.
@@
Table 7: Number of samples from each dataset in the training set.
@@
Table 9: Number of samples from each question type in the training set.
@@ B.1.2 Test Set
The test set comprises 6,122 samples. Tables 10 to 13 provide the sample counts corresponding to each LLM, dataset, prompt template, and question type. Note that datasets with names containing "_enh" refer to the augmented multiple choice question datasets.
@@
Table 10: Number of samples from each LLM in the test set.
@@
Table 11: Number of samples from each dataset in the test set.
@@
Table 12: Number of samples from each prompt template in the test set.
@@
Table 13: Number of samples from each question type in the test set.
@@ B.1.3 Generalization Set
The generalization set comprises 6,468 samples. Tables 14 to 17 provide the sample counts corresponding to each LLM, dataset, prompt template, and question type. Note that datasets with names containing "_enh" refer to the augmented multiple choice question datasets.
@@
Table 14: Number of samples from each LLM in the generalization set.
@@ B.3 Examples from the VAR Dataset
In Figure 4, we provide several representative examples from the VAR dataset. These samples illustrate the structure of the data instances.